OpenStudy (anonymous):

given the following recursive formula write the explicit formula for the sequence t1=2 tn=1/2tn-1 Please help!!

myininaya (myininaya):

\[t_n=\frac{1}{2}t_n-1 ?\]

myininaya (myininaya):

I don't think it can be that...

myininaya (myininaya):

because all you have to do is solve that equation for t_n

myininaya (myininaya):

did you mean \[t_n=\frac{1}{2}t_{n-1}\]?

OpenStudy (anonymous):

it is \[t_n=2*\left( \frac{ 1 }{ 2 } \right)^{n-1}\] so using this \[t_1=2*\left( \frac{ 1 }{ 2 } \right)^{1-1}=2\] and so on \[t_2=2*\left( \frac{ 1 }{ 2 } \right)^{2-1}=1\]